# Low-Loss Space in Neural Networks is Continuous and Fully Connected

### v1

@@ 5 Geometric and Generalization Properties of the Low-Loss Space
We attribute the concentration of high-frequency changes in the first layer to its proximity to the input data, making it more sensitive to high-frequency features that vary significantly across training iterations. This phenomenon is not observed in ResNet18 (Appendix A.3, Figure 7), likely due to residual connections that propagate high-frequency information to deeper layers. This supports the argument that residual connections contribute to flatter loss landscapes and improved generalization [3] by mitigating feature congestion in early layers. Therefore, PCA visualizations of the low-loss space has the potential to serve as a layer-wise diagnostic tool to analyze the types of features captured by each layer.
@@
Figure 4 raises an important question: what is the dimensionality of the low-loss space? Given the symmetric nature of the parameter space [32], no parameter is inherently special. This suggests that all parameters contribute to the low-loss space, and the low-loss space should have the same dimensionality as the parameter space. We empirically verify this intuition through the following analysis.
Since the low-loss space for a single layer lies on a variance sphere in $\mathbb{R}^{d_{l_{x}}}$, this defines an upper bound on its dimensionality. To estimate a lower bound, we examine the subspace spanned by multiple low-loss paths originating from the same trained model. Each path consists of a single training iteration, using a larger batch size and more rounds to ensure movement follows true low-loss directions. We apply singular value decomposition (SVD) to compute the dimensionality of the spanned space.
Empirically, we find that the dimensionality always equals either the minimum of the number of paths or the dimensionality of the parameter space. This indicates that even at a single point, the directions along which loss remains low span the full parameter space. We conclude that the low-loss space has the same dimensionality as the parameter space, implying that no parameter is redundant in defining the low-loss space.
@@
Based on the observation in Figure 3 that models on smaller variance spheres tend to generalize better, we further validate this for ResNet18 on CIFAR100 and ImageNet1k. The results are shown in Figure 5. The most notable finding is from ImageNet1k: using the pretrained start-of-the-art ResNet18 weights from PyTorch [33]—which already achieve 69.758% test accuracy with fine-tuned hyperparameters—Algorithm 2 improves generalization further, reaching 70.55% test accuracy—without additional data augmentation. This suggests that current hyperparameter search methods may be incomplete. In particular, optimal hyperparameters might be layer-specific, since normalization layers are not moved toward the origin in Algorithm 2. As layer-wise hyperparameter tuning is often computationally intensive, exploiting the fully connected structure of the low-loss space may offer a more efficient alternative for improving generalization.
@@ 6 Conclusion
This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-5527. In addition, this work was performed with the support of the Eureka Xecs project TASTI (grant no. 2022005).
@@ Appendix A Technical Appendices and Supplementary Material
A.1 Empirical Validation of Equation 3.1
Figure 6: Layer-wise variance across 100 independently trained ResNet-18 models on the CIFAR-10 dataset, each initialized with a different random seed. The x-axis denotes the model index, and the y-axis represents the variance of each layer’s parameters on a logarithmic scale. Each curve corresponds to one layer, following the PyTorch naming convention. The results show that layer variances remain consistent across different runs. Batch normalization layers are excluded, as their statistics are highly sensitive to the training batches.
A.2 Proof of Equation 6
@@
A.3 Additional PCA results
@@
A.4 Hyperparameter Table
@@
(A1 L4555short for Algorithm 1 Line 4.)
@@
Phase 1 all 10000 train until loss<0.04 or $r$>200 256 SGD($\eta$=0.001, $\beta$=0) 5e-4 0
@@
ResNet18 @CIFAR100 256 SGD($\beta$=0.9, $\lambda$=5e-4) 50

### v3

@@ 5 Geometric and Generalization Properties of the Low-Loss Space
We attribute the concentration of high-frequency changes in the first layer to its sensitivity to the input-level high-frequency features. This effect is less pronounced in ResNet18 (Appendix A.5, Figure 9), likely due to residual connections that propagate high-frequency information to deeper layers. This supports the argument that residual connections contribute to flatter loss landscapes and improved generalization Visualizing_the_Loss_Landscape_of_Neural_Nets by mitigating feature congestion in early layers. Therefore, PCA visualizations of the low-loss space has the potential to serve as a layer-wise diagnostic tool to analyze the types of features captured by each layer.
@@
Figure 4 raises an important question: what is the dimensionality of the low-loss space? Given the symmetric nature of the parameter space lim2024empiricalimpactneuralparameter , no parameter is inherently special. This suggests that all parameters contribute to the low-loss space, and the low-loss space should have the same dimensionality as the parameter space. We empirically verify this intuition through the following analysis.
The low-loss space for a single layer lies on a variance sphere in $\mathbb{R}^{d_{l_{x}}}$, providing an upper bound on its dimensionality. To estimate a lower bound, we construct multiple low-loss paths from the same trained model—each generated via a single training iteration with large batch size and multiple steps to ensure movement follows true low-loss directions. Applying Singular Value Decomposition (SVD) to the set of resulting parameter vectors, we find that the dimensionality of the spanned subspace consistently equals either the minimum of the number of paths or the parameter dimensionality. This confirms that the low-loss space spans the full parameter space, indicating that no parameter is redundant in defining it.
@@
Based on the observation in Figure 3 that models on smaller variance spheres tend to generalize better, we further validate this for ResNet18 on CIFAR100 and ImageNet1k 5206848 . The results are shown in Figure 5. The most notable finding is from ImageNet1k: using the pretrained start-of-the-art ResNet18 weights from PyTorch pytorchResnet18x2014 —which already achieve 69.758% test accuracy with fine-tuned hyperparameters—Algorithm 2 improves generalization further, reaching 70.55% test accuracy—without additional data augmentation. This suggests that current hyperparameter search methods may be incomplete. In particular, optimal hyperparameters might be layer-specific, since normalization layers are not moved toward the origin in Algorithm 2. As layer-wise hyperparameter tuning is often computationally intensive, exploiting the fully connected structure of the low-loss space may offer a more efficient alternative for improving generalization.
@@ 6 Conclusion
This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-5527 and no. EINF-9661. In addition, this work was performed with the support of the Eureka Xecs project TASTI (grant no. 2022005).
@@ Appendix A Technical Appendices and Supplementary Material
A.1 Empirical Validation of Equation 4
Figure 6: Layer-wise weights variance across 100 independently trained ResNet-18 models on the CIFAR-10 dataset, each initialized with a different random seed. The x-axis denotes the model index, and the y-axis represents the variance of each layer’s parameters on a logarithmic scale. Each curve corresponds to one layer, following the PyTorch naming convention. The results show that layer variances remain consistent across different runs. Batch normalization layers are excluded, as their statistics are highly sensitive to the training batches.
A.2 Empirical Validation of Equation 5
Figure 7: Layer-wise weights mean across 100 independently trained ResNet-18 models on the CIFAR-10 dataset. The figure layout follows that of Figure 6. The results indicate that the mean values of most layers are close to zero.
A.3 Proof of Equation 6
@@
A.4 Applying Algorithm 1 on additional models and datasets
[…]
A.5 Additional PCA results
@@
A.6 Hyperparameter Table and Other Configurations
@@
(A1 L4444short for Algorithm 1 Line 4.)
@@
Phase 1 all 10000 train until loss555the loss value is computed using a rolling average with a window size of 10.<0.04 or $r$>200 256 SGD($\eta$=0.001, $\beta$=0) 5e-4 0
@@
ResNet18 @CIFAR100 [40] 256 SGD($\beta$=0.9, $\lambda$=5e-4) 50
@@
Table 4: Dataset preprocessing and augmentation steps. The name of each step aligns with the corresponding function names used in PyTorch. The resulting samples are used for training low-loss models and in the Train step in Algorithm 1 and Algorithm 2.
Data pre-precessing
(augmentation) steps
Step 1 Step 2 Step 3
MNIST Random-rotate for 5 degrees
Random-crop to 28x28
with padding size 2
Normalize with dataset
mean (0.1307) and standard
derivation (0.3081)
CIFAR10 Random-crop to 32x32 with padding size 4
Random-horizontal-flip
with probability 0.5
Normalize with mean
[0.491,0.482,0.446]
and standard derivation
[0.247,0.243,0.262]
CIFAR100 Random-crop to 32x32 with padding size 4
Random-horizontal-flip
with probability 0.5
Normalize with mean
[0.507,0.487,0.441]
and standard derivation
[0.268,0.257,0.276]
ImageNet1k Random-resized-crop to 224x224
Random-horizontal-flip
with probability 0.5
Normalize with mean
[0.485,0.456,0.406]
and standard derivation
[0.229,0.224,0.225]
Table 5: Computation time required to execute Algorithm 1 and Algorithm 2 on our system.
Computation time / simulation Algorithm 1 GPU Algorithm 2 GPU
LeNet5 @MNIST 2 hours Nvidia RTX 3090 20 minutes Nvidia RTX 3090
ResNet18 @CIFAR10 24 hours Nvidia H100 10 hours Nvidia RTX 3090
CCT7_3x1_32 @CIFAR10 48 hours Nvidia H100 24 hours Nvidia RTX 3090
ResNet18 @CIFAR100 4 hours Nvidia RTX 3090
ResNet18 @ImageNet1k 340 hours Nvidia RTX 3090
A.7 Finding Low-Loss Paths with 0.006 threshold for Resnet18@CIFAR10
In this section, we present another hyperparameter setting that could find a low-loss path within $S_{L\leq 0.006}$ ($L_{\text{thres}}=0.006$) for ResNet18 on the CIFAR10 dataset. This threshold is lower than the minimum loss achieved by the low-loss path constructed using the AutoNEB algorithm, which reported a loss of 0.016 for ResNet20 on the CIFAR10 dataset, as shown in Table B.1 of [8]. This result demonstrates that our algorithms can identify low-loss paths with substantially lower loss values.
Figure 10: Training loss (left), layer-wise L2 distance (middle), and testing loss/accuracy (right) of the low-loss path constructed with a threshold of $L_{\text{thres}}=0.006$.
Table 6: Hyperparameters used to find the low-loss path in $S_{L\leq 0.006}$.
Applied layers
Iteration T
(A1 L4)
Train round $r$
(A1 L7)
Batch size $\mathcal{B}$
(A1 L7)
Optimizer and
hyperparameter x
(A1 L7)
step_a
(A1 L14)
step_f
(A1 L14)
ResNet18
@CIFAR10
Phase 1
conv1.weight
bn1.{weigt+bias}
1000 train until $\leq$0.005 or $r$>20000 128 SGD($\eta$=0.001, $\beta$=0) 2e-3 0
Phase 2
Phase 1
layer1.*
1000 2e-3 0
Phase 3
Phase 2
layer2.*
1000 2e-3 0
Phase 4
Phase 3
layer3.*
1000 2e-3 0
Phase 5
Phase 4
layer4.*
1000 2e-3 0
Phase 6
Phase 5
fc.{weight+bias}
3000 2e-3 0
A.8 Modification to Models
To adapt ResNet18 to the image resolution of the CIFAR dataset, the first convolutional layer is modified to use a $3\times 3$ kernel with stride 1 and padding 1. The implementations of ShuffleNet and MobileNet for the CIFAR10 dataset are obtained from [42, 43]. For all other cases, the architecture remains identical to the original implementation.